If 5 samples from 5 populations are given, then what is the probability of getting at least one Type 1 error at 90% confidence interval?


1
Expert's answer
2021-06-09T06:56:52-0400

This is a binomial distribution with p=0.1, n=5.

"P(X\ge1)=1-P(X=0)=1-(1-0.1)^5=0.4095."